Find the sum of the first 5 terms of the sequence defined by an = -7(-4)n − 1 for n ≥ 1.
Answer by stanbon(75887)

Find the sum of the first 5 terms of the sequence defined by an = -7(-4)n − 1 for n ≥ 1.
-------
a(1) = -7(-4)1-1 = 27
a(2) = -7(-4)2-1 = 55
a(3) = -7(-4)3-1 = 83
a(4) = -7(-4)4-1 = 111
a(5) = 139
--------
Sum = 415
================
Cheers,
Stan H.
================